# IKLink: End-Effector Trajectory Tracking with Minimal Reconfigurations

### v1

@@ I INTRODUCTION
The central contribution of this paper is IKLink (III-C, IV). In addition, we describe two efficient but sub-optimal approaches (III-B) as baselines. These three approaches are thoroughly evaluated in a simulation experiment (V) that involves 4 robots and 70 randomly generated end-effector trajectories. Our results show that IKLink generates accurate and smooth motions with fewer reconfigurations when compared to alternative approaches. Finally, we conclude this paper with a discussion of the limitations and implications of this work (VI).
@@ III-B Baseline Approaches
In order to facilitate understanding of the problem, we describe two baseline approaches. These two approaches serve as baselines in our experiment in V. Fig. III-B provides an illustration of the two baseline approaches as well as the method proposed in this paper.
@@
MultiGIK — This approach involves the instantiation of multiple greedy IK solvers in parallel and keeps the longest solution, i.e., the solution that can track the given end-effector trajectory for the maximum duration. This procedure is repeated starting from the subsequent waypoint of the longest solution. MultiGIK has good, but not optimal, performance. Yang et al. [4] have proven that choosing longest possible motion segments leads to a solution with the minimum number of reconfigurations. However, the nested greedy IK solver in MultiGIK is myopic and not guaranteed to find the longest possible motion, so MultiGIK is not optimal. MultiGIK finds optimal solutions only when its greedy motion coincides with the longest possible motion. In practice, MultiGIK works well, so we use it as a baseline in our experiment in V. Our results show that IKLink consistently finds motions with equal or fewer reconfigurations than MultiGIK.
@@ IV TECHNICAL DETAILS
In III-C, we gave an abstract overview of IKLink. We elaborate on the details of our method in this section.
@@ IV-B Dynamic Programming
2
3
[…]
5 for $x=n{-}1$ to 00 do Append ($t[x]$, $S[x,y_{min}]$) to traj $y_{min}\leftarrow p[x,y_{min}]$
return Reverse(traj)
@@ V Evaluation
In this section, we compare our method, IKLink, with two alternative approaches described in III-B, GreedyIK and MutliGIK, on four benchmark tasks. In our prototype systems, MutliGIK instantiates 300 IK solvers in parallel and IKLink samples 300 IK solutions for each waypoint on the reference trajectory ($m=300$). To more fairly assess IKLink over the greedy methods, we provide an additional baseline that takes a similar amount of time to IKLink. The MultiGIK×$\times$30 baseline uses the MultiGIK method but instantiates 300$\times$30 IK solvers in parallel (rather than 300).
@@ V-A Implementation Details
Benchmark Method Mean Num of Reconfig Mean Comput- ation Time (${}^{\dagger}$ Output Joint Motion Metrics Input End-Effector Trajectory Metrics
Mean Joint Vel. (rad/s) Max Pos. Error (${}^{\ddagger}$ Max Rot. Error (${}^{\ddagger}$ Mean # of Waypoints Mean Len. (m) Mean Rot. (rad)
Random Trajectory Tracking
iiwa GreedyIK 13.80$\pm$6.87 0.13$\pm$0.03 0.17$\pm$0.04 1e-3$\pm$6e-6 1e-2$\pm$9e-6 685.6 $\pm$128.2 2.28 $\pm$0.43 9.91 $\pm$2.50
@@
$\ddagger$: In our prototype system, we set the positional and rotational tolerance of IK solvers to be 1e-3 m and 1e-2 rad, respectively. As discussed in IV-A, the accuracy can be improved by allowing more computing time for greedy propagation.
@@ V-E Real-Robot Demonstration
To further demonstrate the effectiveness of IKLink, we generate a welding motion using IKLink and execute it on a physical Rethink Robotics Sawyer robot. The robot is able to accurately and smoothly track the welding trajectory with one reconfiguration. We use $*{}^{*}$ [25] to plan the motion for the reconfiguration. The demonstration is shown in Fig. LABEL:fig:_teaser and the supplementary video [^5] https://youtu.be/EB4bJ6rJtnY.

### v2

@@ I INTRODUCTION
The central contribution of this paper is IKLink (section III-C, section IV). In addition, we describe two efficient but sub-optimal approaches (section III-B) as baselines. These three approaches are thoroughly evaluated in a simulation experiment (section V) that involves 4 robots and 70 randomly generated end-effector trajectories. Our results show that IKLink generates accurate and smooth motions with fewer reconfigurations when compared to alternative approaches. Finally, we conclude this paper with a discussion of the limitations and implications of this work (section VI).
@@ III-B Baseline Approaches
In order to facilitate understanding of the problem, we describe two baseline approaches. These two approaches serve as baselines in our experiment in section V. Fig. III-B provides an illustration of the two baseline approaches as well as the method proposed in this paper.
@@
MultiGIK — This approach involves the instantiation of multiple greedy IK solvers in parallel and keeps the longest solution, i.e., the solution that can track the given end-effector trajectory for the maximum duration. This procedure is repeated starting from the subsequent waypoint of the longest solution. MultiGIK has good, but not optimal, performance. Yang et al. [4] have proven that choosing longest possible motion segments leads to a solution with the minimum number of reconfigurations. However, the nested greedy IK solver in MultiGIK is myopic and not guaranteed to find the longest possible motion, so MultiGIK is not optimal. MultiGIK finds optimal solutions only when its greedy motion coincides with the longest possible motion. In practice, MultiGIK works well, so we use it as a baseline in our experiment in section V. Our results show that IKLink consistently finds motions with equal or fewer reconfigurations than MultiGIK.
@@ IV TECHNICAL DETAILS
In section III-C, we gave an abstract overview of IKLink. We elaborate on the details of our method in this section.
@@ IV-B Dynamic Programming
2     $\Delta t\leftarrow t[x]{-}t[x{-}1]$
3     for $y_{1}=0$ to $m{-}1$ do
4          $c[x,y_{1}],l[x,y_{1}]\leftarrow\infty,\infty$
5          for $y_{2}=0$ to $m{-}1$ do
6               if $c[x{-}1,y_{2}]{+}1{<}c[x,y_{1}]$ or $(c[x{-}1,y_{2}]{+}1$ ${==}c[x,y_{1}]$ and $l[x{-}1,y_{2}]{<}l[x,y_{1}])$ then
[…]
7                    $c[x,y_{1}]\leftarrow c[x{-}1,y_{2}]{+}1$
8                    $p[x,y_{1}],l[x,y_{1}]\leftarrow y_{2},l[x{-}1,y_{2}]$
9
[…]
11                    $d\leftarrow l[x{-}1,y_{2}]{+}\lVert S[x{-}1,y_{2}]-S[x,y_{1}]\rVert_{2}$
12                    if $c[x{-}1,y_{2}]{<}c[x,y_{1}]$ or $(c[x{-}1,y_{2}]{==}$ $c[x,y_{1}]$ and $\,d{<}l[x,y_{1}])$ then
13                         $c[x,y_{1}]\leftarrow c[x{-}1,y_{2}]$
14                         $p[x,y_{1}],l[x,y_{1}]\leftarrow y_{2},d$
15
16
17
18
19
20
21
@@ 25          $c_{min},l_{min},y_{min}\leftarrow c[n{-}1,y],l[n{-}1,y],y$
29 for $x=n{-}1$ to 00 do
30     Append ($t[x]$, $S[x,y_{min}]$) to traj
31     $y_{min}\leftarrow p[x,y_{min}]$
32
33
34return Reverse(traj)
@@ V Evaluation
In this section, we compare our method, IKLink, with two alternative approaches described in section III-B, GreedyIK and MutliGIK, on four benchmark tasks. In our prototype systems, MutliGIK instantiates 300 IK solvers in parallel and IKLink samples 300 IK solutions for each waypoint on the reference trajectory ($m=300$). To more fairly assess IKLink over the greedy methods, we provide an additional baseline that takes a similar amount of time to IKLink. The MultiGIK×$\times$30 baseline uses the MultiGIK method but instantiates 300$\times$30 IK solvers in parallel (rather than 300).
@@ V-A Implementation Details
Benchmark Method Mean Num of Reconfig Mean Comput- ation Time (s) Output Joint Motion Metrics Input End-Effector Trajectory Metrics
Mean Joint Vel. (rad/s) Max Pos. Error (m) Max Rot. Error (rad) Mean # of Waypoints Mean Len. (m) Mean Rot. (rad)
Random Trajectory Tracking iiwa GreedyIK 13.80$\pm$6.87 0.13$\pm$0.03 0.17$\pm$0.04 1e-3$\pm$6e-6 1e-2$\pm$9e-6 685.6 $\pm$128.2 2.28 $\pm$0.43 9.91 $\pm$2.50
@@
$\ddagger$: In our prototype system, we set the positional and rotational tolerance of IK solvers to be 1e-3 m and 1e-2 rad, respectively. As discussed in section IV-A, the accuracy can be improved by allowing more computing time for greedy propagation.
@@ V-E Real-Robot Demonstration
To further demonstrate the effectiveness of IKLink, we generate a welding motion using IKLink and execute it on a physical Rethink Robotics Sawyer robot. The robot is able to accurately and smoothly track the welding trajectory with one reconfiguration. We use RRT [25] to plan the motion for the reconfiguration. The demonstration is shown in Fig. LABEL:fig:_teaser and the supplementary video [^5] https://youtu.be/EB4bJ6rJtnY.